Mountain Gears Inc. has 30% debt, 60% common equity and 10% preferred stock in its capital structure. If Rd=6%, Rp=9% and Re=12%, what is Mountain Gears' cost of capital? Assume the tax rate is 40%.

Group of answer choices

9.18%

7.12%

10.8%

5.27%
Answer #1

After-tax cost of debt=6*(1-tax rate)

=6*(1-0.4)=3.6%

Cost of capital=Respective costs*Respective weight

=(0.3*3.6)+(0.6*12)+(0.1*9)

=9.18%